OpenStudy (anonymous):

A square storage bin that measures 15 feet on a side is filled with corn.If 1 bushel fits in 1.25 cubic feet, how many bushels does it contain? 12 180 2700 3375

OpenStudy (igreen):

Formula: \(V = s^3\) Plug in our side length, 15: \(V = 15^3\) Solve: \(V = 3375\) Divide by 1.25: \(3375 \div 1.25 = 2700\)

OpenStudy (igreen):

So your answer is C, 2700.